To make 2.000 gal of a powdered sports drink, a group of students measure out 2.000 gal of water with 500.-mL, 50.-mL, and 5-mL graduated cylinders. Show how they could get closest to 2.000 gal of water, using these cylinders the fewest times.
1
Expert's answer
2021-09-10T00:42:02-0400

1 gal =

3785.412 ml

2 × 3785.412 = 7570.824ml

15 × 500 = 7500

7570.824-7500 = 70.824

1 × 50 = 50

70.824-50 = 20.824

5 × 4 = 20

20.824 -20 = 0.824ml

= 15 (500ml), 1(50ml), 4(5ml)